In a random sample of 660 full time workers, 132 indicated that they were so angered by a coworker during the past year that they felt like hitting the coworker(but didn't). Compute and interpret a 99% confidence interval for the true proportion of workers so angered in the last year that they wanted to hit a colleague.
Answer #1

n = Sample Size = 660

p = Sample Proportion = 132/660= 0.2

So,

q = 1 - p = 0.8

SE =

= 0.01

ndf = n - 1= 660 - 1 = 659

From Table, critical values of t = 2.5833

Confidence interval:

0.2 (2.5833 X 0.0156)

= 0.2 0.0402

= ( 0.1598 ,0.2402)

Confidence Interval:

0.1598 < P < 0.2402